### v1

# Deep Reasoning Translation via Reinforcement Learning

Abstract
Recently, deep reasoning LLMs (e.g., OpenAI o1/o3 and DeepSeek-R1) have shown promising performance in various complex tasks. Free translation is an important and interesting task in the multilingual world, which requires going beyond word-for-word translation and taking cultural differences into account. This task is still under-explored in deep reasoning LLMs. In this paper, we introduce DeepTrans, a deep reasoning translation model that learns free translation via reinforcement learning. Specifically, we carefully build a reward model with pre-defined scoring criteria on both the translation results and the thought process. Given the source sentences, the reward model teaches the deep translation model how to think and free-translate them during reinforcement learning. In this way, training DeepTrans does not need any labeled translations, avoiding the human-intensive annotation or resource-intensive data synthesis. Experimental results show the effectiveness of DeepTrans. Using Qwen2.5-7B as the backbone, DeepTrans improves performance by 16.3% in literature translation, and outperforms strong deep reasoning baselines as well as baselines that are fine-tuned with synthesized data. Moreover, we summarize the failures and interesting findings during our RL exploration. We hope this work could inspire other researchers in free translation. [^1] https://github.com/krystalan/DRT
Deep Reasoning Translation via Reinforcement Learning
Jiaan Wang,  Fandong Mengthanks:   Corresponding author.,  Jie Zhou Pattern Recognition Center, WeChat AI, Tencent Inc {torchwang,fandongmeng,withtomzhou}@tencent.com
1 Introduction
In recent years, deep reasoning LLMs OpenAI (2024b); Guo et al. (2025) have shown remarkable performance in tasks ranging from math Chen et al. (2025b); Li et al. (2025), codes Zhang et al. (2024), question-taking Guan et al. (2025), etc.
Some researchers bring the success of deep reasoning LLMs to neural machine translation (MT). Marco-o1 Zhao et al. (2024) is introduced for open-ended text generation, and it uses illustrative examples to show the effectiveness of long chain-of-thought (CoT) in colloquial and slang MT. Wang et al. (2024) further systematically study the long CoT in literature MT. They find the literature sentences with metaphors or similes generally require cultural background to understand. Based on this insight, Wang et al. (2024) propose DRT(-o1) LLMs to translate the literature text from English to Chinese with step-by-step reasoning. The DRT LLMs are trained from synthesized long CoT translation data. Chen et al. (2025a) conduct empirical studies on how deep reasoning LLMs work on MT. They verify the strengths of deep reasoning LLMs in historical and cultural translation, but also point out their rambling issues, i.e., LLMs do not follow the MT instruction and fail to translate. Liu et al. (2025) also briefly show the cases of deep reasoning LLMs in MT. More recently, R1-T1 He et al. (2025) is presented, which is the first attempt to employ reinforcement learning (RL) in deep reasoning MT models. Specifically, it uses COMET Rei et al. (2020) score as the reward signal to optimize MT LLMs via modified REINFORCE++ (a RL training algorithm) Hu (2025).
Meanwhile, RL has been demonstrated to have a strong ability to create deep reasoning LLMs Guo et al. (2025); Yu et al. (2025). LLMs can be equipped with powerful reasoning ability solely based on RL training, and such an ability could reveal promising performance on complex tasks (such as math and coding problems). However, RL is still under-explored in deep reasoning MT LLMs. Marco-o1 and DRT are solely trained via supervised fine-tuning (SFT), which requires high computation resources to synthesize data. Though R1-T1 leverages RL during its creation, using the COMET score as the reward signals is not well calibrated. As pointed by Liu et al. (2025), DeepSeek-R1 (a deep reasoning LLM) achieves better translation but a lower COMET compared with DeepSeek-V3 (a non-reasoning LLM). Besides, in some domains, COMET loses its effectiveness in MT evaluation and shows a poor correlation with humans Karpinska and Iyyer (2023).
In this paper, our research goal is to explore how to improve the free translation ability of deep reasoning LLMs via reinforcement learning. Given the strong ability of GRPO Shao et al. (2024), we adopt it as the RL training framework. Consequently, the core problem is how to employ an effective reward model during RL training. There might be three mainstream types of reward models in previous LLM work: i) using widely-adopted MT metrics such as BLEU Papineni et al. (2002), BLEURT Sellam et al. (2020), COMET Rei et al. (2020) and CometKiwi Rei et al. (2022); ii) training a reward model with preference data Stiennon et al. (2020); Nakano et al. (2021), and use the reward model to inference a scalar reward during RL training; iii) designing rule-based rewards Guo et al. (2025). However, these reward models might lose their effectiveness or require high-quality annotated data in MT, hindering their usage. In detail, for i), the reference-based BLEU, BLEURT and COMET are unsuitable for literature MT since the high-quality translation data is extremely difficult to collect. Though CometKiwi is a reference-free metric, its effectiveness is limited in the literature domain, achieving low correlation with humans Wang et al. (2024). For ii), training a reward model requires large-scale annotated preference data, which is also difficult to collect due to the expensive annotation costs. For iii), the rule-based rewards are suitable for tasks whose answers are easy to verify, such as math and code problems Swamy et al. (2025). In MT, we cannot design simple rules to judge the quality of translations.
In view of the strong ability of LLM-as-a-judge Li et al. (2024); Wang et al. (2023); Kocmi and Federmann (2023), we decide to use an advanced instruct LLM as the reward model. Specifically, we carefully design pre-defined scoring criteria on both the translation results and the thought process. For the generation results of deep reasoning MT LLMs, the reward model takes the pre-defined criteria into account to provide a discrete reward (e.g., 3-point or 100-point scores) for RL training. In this manner, we avoid collecting human-annotated preference data for training the reward model. In addition, the effectiveness of the reward model could also be ensured when we adopt the state-of-the-art LLM. We conduct extensive experiments on literature MT, using the above RL strategy to train our DeepTrans model (with the backbone of Qwen2.5-7B). Experimental results show the effectiveness of our method, DeepTrans improves performance by 16.3% and outperforms strong deep reasoning baselines (e.g., QwQ-32B-preview and Marco-o1). Besides, only training with source sentences, DeepTrans outperforms DRT-7B (which is trained on synthesized long thought and translation data) in terms of both automatic metrics and GPT-4o evaluation. Moreover, we share the failures in our pilot RL experiments, and summarize the critical component in reward modeling. We hope our experiments could inspire other researchers in free translation.
Our main contributions are concluded as follows:
  • We propose DeepTrans, which aims at enhancing the free translation ability of deep reasoning LLMs via reinforcement learning. To this end, we use an advanced LLM as the reward model, and carefully design scoring criteria on both translations and thought process.
  • Experimental results on literature translation verify the effectiveness of our DeepTrans. Only training with the source sentences in the RL framework, DeepTrans outperforms strong deep reasoning baselines and LLMs that are fine-tuned with synthesized literature MT data.
  • We summarize the failures and critical components during the RL training to provide a deeper understanding of deep reasoning MT LLMs.
Figure 1: The overview of DeepTrans RL training.
2 DeepTrans
In this section, we introduce DeepTrans. As illustrated in Figure 1, there are three rewards during RL training: format reward, thought reward and translation reward. The first one is obtained via both rules and LLMs, while the latter two rewards are judged by LLMs. We first discuss the rewards designed in the RL framework (§ 2.1) and then provide the training details of DeepTrans (§ 2.2).
2.1 Reward Modeling
Given a source sentence, DeepTrans first thinks and analyzes how to translate the sentence, and finally provides the translation result. We design the format of model generation as “<think> [thought] </think> [translation]”, where “<think>” and “</think>” are two special tokens to indicate the boundary of thought content. “[thought]” and “[translation]” denote the content of thought and translation, respectively. Based on the model generation, we design the following rewards:
Format Reward: We use a regular expression to judge whether the generation format is correct. Besides, in pilot experiments, we find that there might be some explanations in the translation results (i.e., [translation]). To avoid it, we employ DeepSeek-v3 Liu et al. (2024) to judge whether the translation results only contain translations. The judgment prompt is shown as follows:
A translation question requires translating a given text from [src lang] into [trg lang]. The given text is as follows: <text> {src} </text> Someone did this translation task and the translation result is as follows: <translation> {trans} </translation> Please judge whether the translation result belongs to the following situations: 1. It contains only the translation result. 2. It contains the translation result and the explanation. 3. It does not contain the translation result, but only the explanation. Please directly output your judgment result, such as: “Judgment result: 1”, “Judgment result: 2” or “Judgment result: 3”
where “{src}” and “{trans}” denote the source sentence and the translation result, respectively.
If both (1) the generation format is correct, determined by the regular expression, and (2) the translation result only contains translation determined by DeepSeek-v3, we regard the format as correct; otherwise, it is incorrect:
$r_{\text{format}}=\begin{cases}1&\text{if format is correct}\\ 0&\text{if format is incorrect}\end{cases}$ (1)
Thought Reward: The thought process is important to guide the final translation. To reward logical thought processes, we adopt DeepSeek-v3 to provide the thought reward with a 3-point scale.
A translation question requires translating a given text from [src lang] into [trg lang]. The given text is as follows: <text> {src} </text> Someone did this translation question, and began to think how to translate: <think> {think} </think> Please judge whether there is a detailed analysis of the given text in this thinking process: 1. No analysis: Only very shallow thinking was done, and no detailed analysis of the given text was carried out. 2. Slight analysis: The given text was analyzed in detail, and how to translate it was discussed in detail. 3. Detailed analysis: The given text was analyzed in detail, and various translation possibilities were discussed in detail, and trade-offs were made. Please directly output your judgment results, such as: “no analysis”, “slight analysis” or “detailed analysis”
where “{think}” denotes the thought content. The thought reward can be further defined as follows:
$r_{\text{thought}}=\begin{cases}2&\text{if}\operatorname{DS-v3}^{\text{th}}(\text{src},\text{think})=\text{detailed analysis}\\ 1&\text{if}\operatorname{DS-v3}^{\text{th}}(\text{src},\text{think})=\text{slight analysis}\\ 0&\text{if}\operatorname{DS-v3}^{\text{th}}(\text{src},\text{think})=\text{no analysis}\\ \end{cases}$ (2)
where $\operatorname{DS-v3}^{\text{th}}$ denotes using DeepSeek-v3 as the thought reward scorer.
Translation Reward: We also employ DeepSeek-v3 to assess the translation quality. We borrow the prompt from Wang et al. (2024), which is designed to evaluate the quality of literature translation:
SYSTEM PROMPT: Please evaluate the following Chinese translation of an English text. Rate the translation on a scale of 0 to 100, where: - 10 points: Poor translation; the text is somewhat understandable but contains significant errors and awkward phrasing that greatly hinder comprehension for a Chinese reader. - 30 points: Fair translation; the text conveys the basic meaning but lacks fluency and contains several awkward phrases or inaccuracies, making it challenging for a Chinese reader to fully grasp the intended message. - 50 points: Good translation; the text is mostly fluent and conveys the original meaning well, but may have minor awkwardness or slight inaccuracies that could confuse a Chinese reader. - 70 points: Very good translation; the text is smooth and natural, effectively conveying the intended meaning, but may still have minor issues that could slightly affect understanding for a Chinese reader. - 90 points: Excellent translation; the text is fluent and natural, conveying the original meaning clearly and effectively, with no significant issues that would hinder understanding for a Chinese reader. Please provide the reason first, followed by a score. Format your evaluation in the JSON structure below: {"reason": "reason for the score", "score": int} USER PROMPT: <text> {src} </text> <translation> {trans} </translation>
In this manner, DeepSeek-v3 will generate its judgment for a translation and provide its reward with a 100-point scale:
$r_{\text{trans}}=\operatorname{DS-v3}^{\text{tr}}(\text{src},\text{trans})$ (3)
where $\operatorname{DS-v3}^{\text{tr}}$ denotes using DeepSeek-v3 as the translation reward scorer.
Overall Reward: Given the above three types of rewards, we finally design the overall reward:
$r_{\text{all}}=\begin{cases}0&\text{if }r_{\text{format}}=0\\ r_{\text{trans}}+\alpha\times r_{\text{thought}}&\text{if }r_{\text{format}}\neq 0\end{cases}$ (4)
where $\alpha$ is a trade-off hyperparameter between $r_{\text{trans}}$ and $r_{\text{thought}}$ ($\alpha$ is set to 20 in experiments).
2.2 Training Details
Cold Start SFT. We use a non-reasoning LLM, i.e., Qwen2.5-7B-Instruct Yang et al. (2024), as the backbone of DeepTrans. To adapt DeepTrans to the deep reasoning MT task, we first use DeepSeek-R1 Guo et al. (2025) to generate seed translation samples in the general domain (instead of the literature domain). The seed translation samples follow our designed format, and are further used to few-shot SFT the backbone model (named cold start SFT). It is also worth noting that the goal of cold start SFT is not to teach the model free translation but the general translation with long thought.
RL Training. In view of the strong ability of GRPO Shao et al. (2024), which discards the critic model in PPO Schulman et al. (2017). For the policy model $\pi$, given it a source sentence $s$, GRPO first samples a number of generations $\{g_{1},g_{2},...,g_{n}\}$ based on $\pi$, where each $g_{i}$ involves a thought process and the corresponding translation result. Then, GRPO optimizes the policy model $\pi^{\prime}$ by maximizing the following objective:
$\frac{1}{n}\sum_{1}^{n}(\operatorname{min}(\nabla_{\pi}A_{i},\operatorname{clip}(\nabla_{\pi},1-\epsilon,1+\epsilon)A_{i})-\beta\mathcal{D}$ (5)
$\nabla_{\pi}=\frac{\pi^{\prime}(g_{i}|s)}{\pi(g_{i}|s)}$ (6)
$\mathcal{D}=\mathbb{D}_{kl}(\pi^{\prime}||\pi_{ref}))$ (7)
where $\epsilon$ and $\beta$ are hyperparameters. $\pi_{ref}$ is the reference model, and $\mathbb{D}_{kl}(\pi^{\prime}||\pi_{ref}))$ indicates the KL divergence between $\pi^{\prime}$ and $\pi_{ref}$. $A_{i}$ denotes the advantage that is calculated as follows:
$A_{i}=\frac{r_{\text{all}}^{i}-{\operatorname{mean}(\{r_{\text{all}}^{1},r_{\text{all}}^{2},\cdots,r_{\text{all}}^{n}\})}}{{\operatorname{std}(\{r_{\text{all}}^{1},r_{\text{all}}^{2},\cdots,r_{\text{all}}^{n}\})}}.$ (8)
where $r_{\text{all}}^{i}$ denotes the overall reward of $g_{i}$.
3 Experiments
3.1 Experimental Setups
Dataset. Following Wang et al. (2024), we use MetaphorTrans [^2] https://github.com/krystalan/DRT in our experiments, which is a English-to-Chinese literature translation dataset involving 19K training samples and 2K testing samples. Each sample involves a source English sentence, the corresponding thought and Chinese translation. The source sentences are selected from English literature books, and generally contain metaphors or similes, and thus tend to require deep thinking during translation. The thought and translations are synthesized by Qwen2.5-72B. In our experiments, we only use the source sentences of its training samples in our RL training framework.
In addition to the MetaphorTrans test set, we purchase electronic copies of two complete literature books, and use them to evaluate deep reasoning MT models: (1) The Essential O. Henry Collection (by O. Henry) and (2) Orbital (by Samantha Harvey). [^3] Due to copyright protection, we will only release the test data of The Essential O. Henry Collection and not release the content of Orbital (published in 2023).
Metrics. As mentioned by Wang et al. (2024), the references in MetaphorTrans test set are synthesized via LLMs, and are not ensured by human translators. Besides, the golden references of The Essential O. Henry Collection and Orbital are missing. Therefore, we adopt reference-free metrics in our experiments. Specifically, we use “CometKiwi” [^4] https://github.com/Unbabel/COMET Rei et al. (2022) to evaluate the model translations, which uses a language model to judge whether a model translation conveys the semantics of the source sentence. Moreover, following Wang et al. (2024), we use evaluators implemented using GPT-4o in two reference-free manners, which we refer to as “GRF” and “GEA”, respectively. The evaluation prompt of “GRF” borrows from Kocmi and Federmann (2023), and are illustrated in Appendix A. For “GEA”, the prompt mainly borrows from Wang et al. (2024), and we employ two variants of “GEA”, i.e., “GEA100” and “GEA5”. The evaluation prompt of “GEA100” is the same as the translation reward illustrated in § 2.1, while that of “GEA5” simply narrows the scoring scope of “GEA100” from a 100-point to a 5-point scale. Among them, “GRF” evaluates translations from a general perspective while “GEA5” and “GEA100” evaluate translation from a literary perspective. Since GRF, GEA5 and GEA100 need the costs of OpenAI’s API, we randomly select 400 samples from each test set (i.e., MetaphorTrans, O. Henry and Orbital), to conduct evaluation.
Backbones. We try to use LLMs with less than 10B parameters as the backbone of DeepTrans, given the high computation costs during RL training. Among all open-source LLMs, Qwen2.5-7B Yang et al. (2024) and Llama3-8B Grattafiori et al. (2024) are state-of-the-art choices. As pointed by Wang et al. (2024), Qwen2.5-7B-Instruct outperforms Llama3-8B-Instruct in literature translation. Therefore, we finally use Qwen2.5-7B-Instruct as the backbone of DeepTrans.
Para. Deep Training on Ckpt.
Reason. MetaphorTrans
General Non-reasoning Baselines
Llama-3.1-8B-Instruct 8B $\times$ $\times$ [Uncaptioned image]
Qwen2.5-7B-Instruct 7B $\times$ $\times$ [Uncaptioned image]
Qwen2.5-14B-Instruct 14B $\times$ $\times$ [Uncaptioned image]
GPT-4o - $\times$ $\times$ -
General Reasoning Baselines
Marco-o1-7B 7B $\times$ [Uncaptioned image]
QwQ-32B-preview 32B $\times$ [Uncaptioned image]
QwQ-32B 32B $\times$ [Uncaptioned image]
DeepSeek-Qwen-7B 7B $\times$ [Uncaptioned image]
DeepSeek-Llama-8B 8B $\times$ [Uncaptioned image]
DeepSeek-Qwen-14B 14B $\times$ [Uncaptioned image]
DeepSeek-Qwen-32B 32B $\times$ [Uncaptioned image]
DeepSeek-R1 671B $\times$ [Uncaptioned image]
o1-preview - $\times$ -
MT Non-reasoning Baselines
Llama-3.1-8B-SFT 8B $\times$ -
Qwen2.5-7B-SFT 7B $\times$ -
Qwen2.5-14B-SFT 14B $\times$ -
MT Non-reasoning Baselines
DRT-7B 7B [Uncaptioned image]
DRT-8B 8B [Uncaptioned image]
DRT-14B 14B [Uncaptioned image]
Our
DeepTrans-7B 7B
Table 1: Comparisons between baselines and DeepTrans. Para.: Parameter; Reason.: Reasoning. Ckpt.: Checkpoint. “Training on MetaphorTrans” indicates whether the model is trained on the MetaphorTrans dataset.
Model MetaphorTrans O. Henry Orbital
GRF GEA5 GEA100 CometKiwi GRF GEA5 GEA100 CometKiwi GRF GEA5 GEA100 CometKiwi
Llama-3.1-8B-Instruct 79.25 3.31 59.58 70.14 79.73 3.43 57.17 74.35 79.92 3.54 59.90 75.09
Qwen2.5-7B-Instruct 81.53 3.62 66.21 70.36 85.26 3.83 66.50 76.18 83.38 4.00 70.10 76.46
Qwen2.5-14B-Instruct 84.74 3.87 70.86 72.01 86.83 3.98 70.53 77.04 84.39 4.09 71.65 76.55
Marco-o1-7B 82.41 3.57 64.24 71.62 83.12 3.71 63.11 76.00 81.84 3.89 67.64 75.38
DeepSeek-Qwen-7B 65.16 2.67 43.66 63.49 68.97 2.86 45.64 70.67 71.28 3.16 51.91 72.43
DeepSeek-Llama-8B 76.31 3.24 56.89 67.13 78.17 3.39 56.14 73.39 78.91 3.64 59.75 74.47
DeepSeek-Qwen-14B 83.92 3.81 70.64 71.01 83.27 3.82 64.79 75.22 82.30 4.01 69.10 76.28
Llama-3.1-8B-SFT 84.10 3.88 69.33 70.25 85.04 3.87 66.60 76.14 80.37 3.87 64.38 75.11
Qwen2.5-7B-SFT 85.06 3.93 72.29 71.03 86.84 4.05 71.05 77.29 85.46 4.12 70.55 76.32
Qwen2.5-14B-SFT 85.66 4.02 74.53 72.08 87.27 4.05 73.06 77.54 85.55 4.14 75.84 77.40
DRT-7B 85.57 4.05 75.05 71.78 86.36 3.96 69.51 76.12 81.69 3.84 65.56 69.95
DRT-8B 84.49 3.91 69.65 70.85 83.61 3.75 64.76 73.89 79.14 3.65 61.36 66.36
DRT-14B 87.19 4.13 77.41 72.11 87.38 4.00 72.59 76.70 82.19 3.98 69.36 70.99
DeepTrans-7B (Our) 88.84 4.21 75.38 71.82 87.95 4.22 76.92 77.04 87.95 4.22 76.92 76.65
Table 2: Comparison results of DeepTrans and open-source baselines (< 30B parameters). The bold and the underline denote the best and second-best performances, respectively. $\dagger$” and“$\ddagger$” denote statistically significant better than the DRT-14B with t-test p < 0.01 and 0.05, respectively. $\diamondsuit$” denotes models are trained on MetaphorTrans.
Model MetaphorTrans O. Henry Orbital
GRF GEA5 GEA100 CometKiwi GRF GEA5 GEA100 CometKiwi GRF GEA5 GEA100 CometKiwi
GPT-4o 85.57 3.86 71.88 73.01 88.30 4.00 71.06 76.74 85.91 4.17 73.54 77.67
o1-preview 87.11 4.06 78.01 73.70 89.73 4.14 76.17 78.41 86.85 4.26 76.80 78.86
QwQ-32B-preview 86.31 4.00 75.50 71.48 87.61 4.03 70.79 76.86 84.79 4.04 71.03 76.17
QwQ-32B 88.06 4.09 74.38 72.88 88.02 4.21 76.36 77.71 87.83 4.15 76.55 77.55
DeepSeek-Qwen-32B 84.78 3.87 71.88 71.93 87.03 4.03 70.81 76.75 85.36 4.16 73.62 77.80
DeepSeek-R1 84.29 4.02 73.78 68.33 89.79 4.17 77.03 77.01 87.37 4.27 80.06 76.17
DeepTrans-7B (Our) 88.84 4.21 75.38 71.82 87.95 4.22 76.92 77.04 87.95 4.22 76.92 76.65
Table 3: Comparison results of DeepTrans and baselines (> 30B parameters or commercial LLMs). The bold and the underline denote the best and second-best performances, respectively.
3.2 Baselines
(1) General Non-reasoning LLMs. We use Llama-3.1-8B-Instruct Grattafiori et al. (2024), Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct Yang et al. (2024) and GPT-4o OpenAI (2024a) as baselines.
(2) General Reasoning LLMs. Marco-o1-7B Zhao et al. (2024), QwQ-32B-preview, QwQ-32B Qwen (2024), DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Qwen-14B, DeepSeek-R1-Distill-Qwen-32B, DeepSeek-R1 Guo et al. (2025) and o1-preview OpenAI (2024b) are used as baselines.
(3) MT Non-reasoning LLMs. Wang et al. (2024) fine-tune three LLMs with only paired sentences of the MetaphorTrans training data (without thought). This setting allows LLMs to learn the mapping from source literature sentences to the corresponding Chinese translations directly. The fine-tuned LLMs are denoted as Llama-3.1-8B-SFT, Qwen2.5-7B-SFT and Qwen2.5-14B-SFT.
(4) MT Reasoning LLMs. Wang et al. (2024) introduce DRT-7B, DRT-8B and DRT-14B models, which are fine-tuned on the whole MetaphorTrans training data (with thought). Given sentences, these LLMs could first think about how to translate them and then provide the corresponding translations.
To make a deeper understanding of these baselines, we summarize their key features in Table 1.
3.3 Implementation Details.
Cold Start SFT. We randomly select 4K English sentences from WMT24555https://www2.statmt.org/wmt24/index.html in the general domain. Then, DeepSeek-R1 is employed to translate these sentences from English to Chinese in a deep reasoning manner. The synthesized 4K samples are further used to SFT DeepTrans (i.e., cold-start SFT). In detail, Llama-Factory framework Zheng et al. (2024) is used during the SFT stage. We conduct experiments on 8$\times$×NVIDIA H20 GPUs (96G) with 1e-5 learning rate and 8 (8$\times$1) batch size. DeepSpeed ZeRO-3 optimization Rasley et al. (2020) is also used during SFT. We set the number of SFT epochs to 2, and it costs about 1 GPU hour.
RL Training. We use GRPO RL algorithm implemented by verl [^6] https://github.com/volcengine/verl. During training, 2$\times$8 H20 GPUs are used, where 8 GPUs are used to deploy DeepSeek-v3 (awq quantization) as the reward model, and another 8 GPUs are used to optimize the policy model. We set the batch size to 64, the learning rate to 1e-6, the rollout number to 8 and the rollout temperature to 0.6, and the KL loss coefficient to 1e-3. The number of training epochs is set to 2. The RL training costs 2K GPU hours.
(a) Changes in overall rewards
(b) Changes in thought length
(c) Changes in CometKiwi
(d) Changes in GRF
(e) Changes in GEA5
(f) Changes in GEA100
Figure 2: Performance changes of DeepTrans-7B during RL training. Among them, (b), (d), (e) and (f) are conducted on the MetaphorTrans test set, while (c) is conducted on all three test sets. The training process totally involves 2 epochs and each epoch involves 300 training steps.
3.4 Main Results
As shown in Table 2, the results demonstrate the effectiveness of DeepTrans-7B. Specifically, it outperforms baselines with less than 7B parameters in terms of all metrics. Compared with all baselines (< 30B parameters), DeepTrans-7B achieves state-of-the-art performance in most metrics, especially GEA5 and GRF. Specifically, DeepTrans-7B outperforms its backbone (Qwen2.5-7B-Instruct) by 16.3%, 13.8% and 9.0% in terms of GEA5, GEA100 and GRF on the MetaphorTrans test set. As for strong baselines that are fine-tuned on MetaphorTrans (denoted with $\diamondsuit$”), they use both the source sentences and the synthesized translation targets. In contrast, DeepTrans-7B only uses the source sentences and achieves promising results. It is worth noting that training with RL does not need the synthesized translations (and thought process), and thus avoids the data quality issue as well as potential bias in synthesized data.
We also compare DeepTrans-7B with baselines (> 30B parameters or commercial LLMs), the results are shown in Table 3. DeepTrans-7B generally outperforms QwQ-32B-preview and DeepSeek-Qwen-32B, and achieves competitive results with GPT-4o, o1-preview, QwQ-32B and DeepSeek-R1.
Source Sentence DeepTrans (Cold Start) DeepTrans (150 Steps) DeepTrans (300 Steps) DeepTrans (450 Steps) DeepTrans (600 Steps)
The man from the West unfolded the little piece of paper handed him. His hand was steady when he began to read, but it trembled a little by the time he had finished. The note was rather short. {CJK}UTF8gbsn 西方人展开递给他的一小张纸条。他开始阅读时手很稳,但读完后微微发抖。那张纸条很短。 [CometKiwi: 80.10] {CJK}UTF8gbsn 首先,西边来的男人展开递给他的一小张纸。他开始阅读时双手稳当,但读完后手指已微微颤抖。那张便条颇为简短。 [CometKiwi: 75.89] {CJK}UTF8gbsn 那位西部来的人接过那张折叠的小纸片,启口之际,双手仍显镇定,及至读罢,指尖却微微颤动。此笺虽属简短,却已道尽千般情愫[CometKiwi: 76.63] {CJK}UTF8gbsn 这位来自西方的男子,缓缓展开了递到他手中的小小纸片,当他开始阅读时,双手依然镇定,然而等到读罢之时,指尖已微微颤动。那张便条颇为简短。 [CometKiwi: 82.32] {CJK}UTF8gbsn 来自西方的男子,缓缓展开了递到他手中的小小纸片,起初,他的手显得十分稳当,可等到读罢,指尖竟微微颤抖起来。这封便条其实颇为简短。 [CometKiwi: 83.42]
When it was light enough Johnsy, the merciless, commanded that the shade be raised. {CJK}UTF8gbsn 当光线足够明亮时,无情的约翰西斯命令把窗帘拉上。 [CometKiwi: 81.72] {CJK}UTF8gbsn 当光线足够明亮时,那个无情的约翰西便下令拉起窗帘。 [CometKiwi: 81.56] {CJK}UTF8gbsn 待天色稍霁,约翰西这颗冷酷的心,便命将帘幕徐徐卷起[CometKiwi: 57.61] {CJK}UTF8gbsn 等到光线足够了,这个冷酷无情的约翰西,便下令将那阴影缓缓卷起。 [CometKiwi: 72.48] {CJK}UTF8gbsn 约翰西,这个冷酷无情的人,等到天色足够明亮,便命令将帘幕缓缓拉起。 [CometKiwi: 77.28]
When one loves one’s Art no service seems too hard. {CJK}UTF8gbsn 当一个人热爱自己的艺术时,任何服务都显得不值一提。 [CometKiwi: 70.98] {CJK}UTF8gbsn 一旦钟爱艺术,便觉任何奉献皆非难事。 [CometKiwi: 81.86] {CJK}UTF8gbsn 一旦钟爱自家的艺术,任是何般劳顿艰辛,也觉微不足道,甘之如饴。 [CometKiwi: 65.53] {CJK}UTF8gbsn 只要钟爱自己的艺术,再大的付出也觉甘愿。 [CometKiwi: 67.99] {CJK}UTF8gbsn 只要钟爱自己的艺术,再大的付出也绝不嫌苦。 [CometKiwi: 76.16]
For two or three days she went out canvassing for pupils. One evening she came home elated. {CJK}UTF8gbsn 连续两三天她出去招生,一个傍晚她兴高采烈地回家了。 [CometKiwi: 83.55] {CJK}UTF8gbsn 连续两三天,她外出物色学生。某晚归来时,她显得欣喜若狂。 [CometKiwi: 84.96] {CJK}UTF8gbsn 接连三四日,她外出寻觅门徒踪迹,直至某晚归来,已得诸多佳音满怀欣喜。 [CometKiwi: 64.59] {CJK}UTF8gbsn 连续一两天,她便出门去物色学生。一天傍晚,她兴冲冲地回到了家中。 [CometKiwi: 82.87] {CJK}UTF8gbsn 连续两三天,她外出寻觅学生上门求教。一天傍晚,她满面春风地回到了家中。 [CometKiwi: 81.21]
Table 4: Case Studies on O. Henry. Purple indicates the translations are in strong Chinese classical literary style.
3.5 Intermediate-Stage Analyses
To provide a deeper analysis of DeepTrans-7B, we further discuss the performance changes during RL training. Figure 2 shows the corresponding details, which we analyze from the following aspects:
In terms of overall rewards, the training rewards and validation rewards generally increase along with the training process. The full score of the overall reward is 140 (c.f. Eq. 4), and DeepTrans-7B finally reaches about 120, which means $r_{\text{trans}}$ reaches at least 80 (c.f. Eq. 3). According to the definition of $r_{\text{trans}}$, 80 points indicate the model translations are between “very good translation” and “excellent translation”, showing the significant improvement brought by RL training. Similarly, as shown in Figure 2 (d), (e) and (f), model performance (in terms of GRF, GEA5 and GEA100) also generally increases along with the training process.
In terms of thought length, the average length first significantly decreases from 800+ to 100 tokens in the first 100 steps, and then slowly increases to 200 tokens. This finding indicates that during RL training, the model will first focus on the translation results, and try to provide translations approved by the reward model. Meanwhile, the model starts to simplify the thought content, since no additional thought rewards are given. However, when the thought length is below a threshold, the model cannot provide a meaningful thought process with short words, and consequently, it will be punished by $r_{\text{thought}}$. Therefore, the model will stop simplifying the thought content, and learn to generate better translations and keep high-quality thought processes simultaneously. Later, during RL, the model will find a deeper thought that will lead to a better translation, and increase its thought length.
In terms of CometKiwi, we find an interesting phenomenon: the performance drops a lot at first, and finally returns to the initial performance. To understand the reason behind this, we provide several cases generated by the intermediate-stage DeepTrans models in Table 4. We find that the model trained with 300 steps has a special characteristic: its translations are in strong Chinese classical literary style (e.g., ancient poetry). We mark several classical-style terms in purple, and these terms are also not commonly used in modern Chinese. The classical style might not be well captured by the CometKiwi calculating model [^7] https://huggingface.co/Unbabel/wmt22-cometkiwi-da, thus resulting in a low CometKiwi score. However, it still can be recognized by the GPT-4o evaluators. As shown in Figure 2, DeepTrans (300 steps) still outforms DeepTrans (cold start) in terms of GRF, GEA5 and GEA100. For this stylization phenomenon, we also want to discuss the following questions:
(1) Is stylization better or not? Though classical-style translations will provide some interesting results, we think this might lose the generalizability of the MT models. The classical-style translations are not suitable for all genres in literature translation. Our goal is to make the model provide translations that should be more easily absorbed by native people in target languages. The classical-style translations do not follow this goal, and it should be considered as a bias or preference. Therefore, we recommend that future work should also evaluate the stylization phenomenon during RL training.
(2) Why does DeepTrans first drop into the stylization phenomenon and finally get out of it? First, the model recognizes that the classical-style translations will receive a high translation reward (i.e., $r_{\text{trans}}$) during RL training, and thus, the model starts to provide classical-style translations. This also indicates the bias/preference of the reward model (i.e., DeepSeek-v3), which thinks classical-style translations are easy to understand for Chinese. However, such a stylization strategy will trap the model into a local optimal solution. Fortunately, the model discovers other better solutions during RL training, and finally discards the stylization strategy. Another important question is naturally raised: which factors lead the model to get out of the local optimal solution? We discuss it in § 4.3.
Figure 3: The performance comparisons between DeepTrans ($r_{\text{all}}$) and DeepTrans ($r_{\text{all (w/ lp)}}$) on MetaphorTrans test set.
(a) In terms of GRF
(b) In terms of CometKiwi
Figure 4: The performance comparisons among models trained with translation rewards at different scoring scales. The results are evaluated on the MetaphorTrans test set.
4 Failure Experience
In this section, we discuss our failures in reward design. These experiences could also be regards as the ablation study in $r_{\text{all}}$.
4.1 The Length Penalty of Thought Process
In the pilot RL experiments, we also design a length penalty for the thought content. Specifically, it could be defined as follows:
$r_{\text{length}}=\begin{cases}-\frac{\operatorname{length}(\text{think})-\beta}{\eta}&\text{if}\operatorname{length}\text{(think)}\geq\beta\\ 0&\text{if}\operatorname{length}\text{(think)}<\beta\end{cases}$ (9)
[…]
$r_{\text{all (w/ lp)}}=\begin{cases}0&\text{if }r_{\text{format}}=0\\ r_{\text{trans}}+\alpha\times(r_{\text{thought}}+r_{\text{length}})&\text{if }r_{\text{format}}\neq 0\end{cases}$ (10)
We set $\beta$ to 400, $\eta$ to 400, and $\alpha$ to 20 in the pilot experiments. The DeepTrans-7B model trained with $r_{\text{all (w/ lp)}}$ is denotes as DeepTrans ($r_{\text{all (w/ lp)}}$). As shown in Figure 3, we find that the length penalty does not bring additional improvement to model performance during training. In contrast, it reduces the rate of improvement in model performance. Therefore, we do not adopt the length penalty in the final overall reward.
4.2 The Effect of Thought Reward
We also attempt to remove the thought reward during RL training, and the variant overall reward is defined as:
$r_{\text{all (w/o th)}}=\begin{cases}0&\text{if }r_{\text{format}}=0\\ r_{\text{trans}}&\text{if }r_{\text{format}}\neq 0\end{cases}$ (11)
However, during RL training, we find that DeepTrans ($r_{\text{all (w/o th)}}$) quickly discards the thought process, and tends to directly output the final translation until the end of the training.
To make the model provide a valuable thought process, we also consider a short length penalty on the thought process. In detail, if the thought length is less than $k$ tokens, we give a huge penalty to the overall reward:
$r_{\text{all (w/o th \& w/ sp)}}=\begin{cases}0&\text{if }r_{\text{format}}=0\\ 0&\text{elif }\operatorname{length}\text{(think)}<10\\ r_{\text{trans}}&\text{else}\end{cases}$ (12)
In this way, the final trained model tends to only provide several empty words in the thought content, e.g., “Well, first I have to understand the meaning of the sentence and make sure the translation is accurate. Then, I will express it in fluent Chinese, keeping the sentence coherent and natural”. This kind of thinking process loses the relevance and specificity towards source sentences.
Besides, the model trained with $r_{\text{all (w/o th)}}$ or $r_{\text{all (w/o th \& w/ sp)}}$ significantly underperforms DeepTrans-7B in terms of all metrics, and also shows code-mix results in the translations. These findings demonstrate the importance of the thought process in model application, and we should consider how to guide models to generate valuable thought during RL training.
Source Sentence DeepTrans (Cold Start) DeepTrans Variants (300 Steps) DeepTrans Variants (600 Steps)
TR. w/ 3-point scale TR. w/ 5-point scale TR. w/ 3-point scale TR. w/ 5-point scale
The man from the West unfolded the little piece of paper handed him. His hand was steady when he began to read, but it trembled a little by the time he had finished. The note was rather short. {CJK}UTF8gbsn 西方人展开递给他的一小张纸条。他开始阅读时手很稳,但读完后微微发抖。那张纸条很短。 [CometKiwi: 80.10] {CJK}UTF8gbsn 那自西而来之人展开了递与他之小小书笺,启阅时手心尚稳,及至终了,已微颤难禁。此简不过寥寥数语[CometKiwi: 66.25] {CJK}UTF8gbsn那自西方而来的男子轻轻展开递至掌心的小小纸笺初时指尖稳健如常,逐字细阅其上墨迹,及至文末,方觉指间微颤不宁。此纸不过寥寥数行,却藏无限意味[CometKiwi: 68.79] {CJK}UTF8gbsn 西边来人展信细阅初启笔端神定如磐,及至文终,指尖微颤。简札不过寥寥数行[CometKiwi: 74.02] {CJK}UTF8gbsn 西边来的人,缓缓展开递到手里的小小纸片。起初,他的双手握得十分稳当,可待读罢之时,指尖已微微有些发颤了。那张纸条实在不算长。 [CometKiwi: 79.41]
When it was light enough Johnsy, the merciless, commanded that the shade be raised. {CJK}UTF8gbsn 当光线足够明亮时,无情的约翰西斯命令把窗帘拉上。 [CometKiwi: 81.72] {CJK}UTF8gbsn 当光线已然充足,约翰西,那无情之人,竟令帷幕徐徐卷起[CometKiwi: 74.46] {CJK}UTF8gbsn 待天色渐明,那冷漠无情的约翰西,便命人将帘幕徐徐卷起[CometKiwi: 75.08] {CJK}UTF8gbsn 天色稍显明亮,约翰西,那颗铁石心肠,便下令将阴影卷起。 [CometKiwi: 66.96] {CJK}UTF8gbsn 待天色足够明亮,约翰西这颗铁石心肠,已然下令将帘幕卷起。 [CometKiwi: 71.40]
When one loves one’s Art no service seems too hard. {CJK}UTF8gbsn 当一个人热爱自己的艺术时,任何服务都显得不值一提。 [CometKiwi: 70.98] {CJK}UTF8gbsn 若钟情于艺道,任劳任怨亦无妨[CometKiwi: 69.38] {CJK}UTF8gbsn 一旦钟爱所钟之艺便觉任劳任怨非艰[CometKiwi: 69.38] {CJK}UTF8gbsn 艺海倾心无难事,任凭千般苦亦甘[CometKiwi: 54.76] {CJK}UTF8gbsn 只要钟爱所钟情的艺术,任何付出都觉轻而易举。 [CometKiwi: 76.59]
For two or three days she went out canvassing for pupils. One evening she came home elated. {CJK}UTF8gbsn 连续两三天她出去招生,一个傍晚她兴高采烈地回家了。 [CometKiwi: 83.55] {CJK}UTF8gbsn 连续几日外出寻觅门徒一夕归家心欢悦[CometKiwi: 80.25] {CJK}UTF8gbsn 她连着几日四处奔走寻觅门徒,直至某晚,她满心欢喜地踏入了家园。 [CometKiwi: 74.54] {CJK}UTF8gbsn 连续几日奔走寻门徒,直至傍晚方归家满心欢喜步履轻,笑颜如花映晚霞[CometKiwi: 55.10] {CJK}UTF8gbsn 她接连走了两三天,到处奔走寻觅学生。一天傍晚,她满面春风地回到了家中。 [CometKiwi: 83.87]
Table 5: Case Studies on O. Henry (TR.: Translation Reward). Purple indicates the translations are in strong Chinese classical literary style.
4.3 The Effect of Translation Reward
We define the translation reward with a 100-point scale in § 2.1, and we further study the effect of the scoring criteria in translation rewards. Specifically, we use two variants based on the original $r_{\text{trans}}$: $r_{\text{trans (3 point)}}$ and $r_{\text{trans (5 point)}}$ simply narrow the scoring scope of $r_{\text{trans}}$ from a 100-point to a 3-point scale and a 5-point scale, respectively. During training with these two variants, the $\alpha$ in Eq. 4 is also changed to ensure the same ratio between thought rewards and translation rewards. We denote the models trained with the variants as DeepTrans (trans reward w/ 3-point scale) and DeepTrans (trans reward w/ 5-point scale), respectively. As shown in Figure 4 (a), both variants show the sub-optimal results in terms of GRF, verifying the rationality of our original design.
As for CometKiwi scores in Figure 4 (b), we find that DeepTrans (trans reward w/ 5-point scale) shows similar trends with original DeepTrans, i.e., first drops and finally recovers. However, DeepTrans (trans reward w/ 3-point scale) continues to decline and fluctuate during the training progress. In the end, it achieves extremely low performance in terms of CometKiwi, and perhaps it needs more training steps to recover the performance. In view of the stylization phenomenon we discussed in § 3.5 and Table 4, we conjecture the trends shown in Figure 4 (b) are attributed to stylization.
To figure it out, we provide case studies on the variant models in Table 5 (where the cases are the same as Table 4). As we can see, when the models are trained with 300 steps, both variants show the stylization phenomenon. Further, when trained with 600 steps, we find that DeepTrans (trans reward w/ 5-point scale) discards the classical-style translations. In contrast, DeepTrans (trans reward w/ 3-point scale) enhances the classical-style translations, and the results are like ancient poetry.
As we also described in § 3.5, the basic reason of the stylization phenomenon is a bias/preference of the reward model, which encourages the translation model to provide stylized translations as a local optimal solution. When the reward model provides a wider scoring scope (e.g., the original $r_{\text{trans}}$ with 100-point scale), it gives a larger latent space for the policy model to explore. Therefore, the policy model is more likely to get out of the local optimal solution. If using a narrow reward scope, e.g., $r_{\text{trans (3 point)}}$, some minor changes in the policy model will not be reflected by the reward signal immediately, and thus make it hard to transfer the translation strategies. In conclusion, the scoring scope of the reward models has a great effect on the policy model, and it is necessary to employ a wider reward scope for the model to explore a good translation strategy.
5 Related Work
5.1 Deep Reasoning LLMs
In recent years, deep reasoning LLMs, e.g., OpenAI o1 OpenAI (2024b) and DeepSeek-R1 Guo et al. (2025), have pioneered a growing research in long chain-of-thought (CoT) reasoning. Many studies make their efforts on different tasks or scenarios Zhang et al. (2024); Chen et al. (2025b); Li et al. (2025); Guan et al. (2025); Jin et al. (2025). Chen et al. (2025b) provide a systematic survey on deep reasoning LLMs to introduce the development process. Some researchers investigate the MT capability of deep reasoning LLMs. Zhao et al. (2024) and Liu et al. (2025) briefly show that the long CoT reasoning helps the model to reach more idiomatic translations. DRT(-o1) Wang et al. (2024) is further proposed in literature MT, and it is trained with synthesized thought processes as well as translations. An empirical study Chen et al. (2025a) also discusses how well deep reasoning LLMs perform on MT tasks. R1-T1 He et al. (2025) leverages COMET scores as reward signals to train deep reasoning MT models with RL. However, the deep reasoning LLMs in MT are still under-explored.
5.2 Literature Translation
Literature translation is more difficult than news translation since it often requires equivalence beyond the word level Thai et al. (2022). Early literature translation work is limited to small-scale attempts and uses human evaluation to judge model performance Genzel et al. (2010); Jones and Irvine (2013); Besacier and Schwartz (2015); Toral et al. (2018). Recently, Karpinska and Iyyer (2023) utilize GPT-3.5 to perform literature translation, and show that the discourse-level information could enhance model performance. Thai et al. (2022) propose the Par3 dataset, which aims to benchmark LLMs’ literature translation capability from non-English languages to English. Wang et al. (2024) contribute the MetaphorTrans dataset with English-to-Chinese translation. Besides, the sentences in MetaphorTrans generally involve metaphors, thus, it might be suitable for deep reasoning LLMs to perform. DRT Wang et al. (2024) is further proposed based on MetaphorTrans by supervised fine-tuning.
5.3 MT via RL
Machine translation via reinforcement learning has been explored in the literature. Early work tries to train MT models via optimizing BLEU scores Ranzato et al. (2016); Shen et al. (2016); Bahdanau et al. (2017). Wu et al. (2016) design GLEU scores as the rewards to deal with the drawbacks of BLEU in single sentence evaluation. Wu et al. (2018) provide an empirical study on MT via RL, and they propose a method to involve large-scale monolingual data during RL training. Choshen et al. point out the challenges of optimizing MT models via RL, e.g., high-dimensional discrete action space and sparse reward signals. Kiegeland and Kreutzer (2021) provide further analyses on these challenges. Kang et al. (2020) study document-level MT, and they propose a new method to both select context and translate sentences via RL. Though great performance has been achieved, most previous work only focuses on non-reasoning models. After the emergence of deep reasoning LLMs, R1-T1 He et al. (2025) is proposed to optimize deep reasoning LLMs via RL, and it uses COMET scores as the reward signals. However, using COMET as the optimization goal is not well calibrated. In the literature domain, COMET shows low effectiveness as an evaluation metric Karpinska and Iyyer (2023).
6 Conclusion
In this work, we study how to improve the free translation ability of deep reasoning LLMs via RL. We design three types of reward signals in RL: format reward, thought reward and translation reward. Among them, rules and DeepSeek-v3 are used to provide discrete reward values. In this way, using Qwen2.5-7B-Instruct as backbone, we train DeepTrans-7B via GRPO RL algorithm. The experimental results in the literature translation show the effectiveness of DeepTrans-7B. It outperforms QwQ-32B-preview and DeepSeek-Qwen-32B LLMs, and achieves competitive results with DeepSeek-R1 and OpenAI’s o1. In addition, we give deeper analyses on model results during RL training, and find a stylization phenomenon that needs to be carefully considered. Moreover, we summarize the failure experiences and critical components during RL training to provide a deeper understanding of deep reasoning MT LLMs.
Appendix A GRF Evaluator
The GRF prompt borrows from Kocmi and Federmann (2023):
Score the following translation from English to Chinese on a continuous scale from 0 to 100, where score of zero means "no meaning preserved" and score of one hundred means "perfect preservation of meaning, with faithfulness, expressiveness, and elegance". English source: {src} Chinese translation: {hyp} Score: